# Infinite graph product of groups I: Geometry of the extension graph

### v2

@@ 3.2 Connection to the crossing graph and the contact graph
$E(\widehat{\Gamma}^{e})$ =$=\{(x,y)\in V(\Gamma^{e})\times V(\Gamma^{e})\mid\text{ $x\neq y$and $\exists\,g\in\Gamma\mathcal{G},\{x,y\}\subset g.\Gamma$}\}$
@@
[…]
@@
Lemma 3.21.
Proof.
$C(q,X)=\{p\in 2^{V(\Gamma^{e})}\mid\text{$p$ is a circuit in $X$such that $|p|=\mathrm{girth}(\Gamma)$ and $q\subset V(p)$}\}.$
@@ 3.3 Admissible paths
$\mathscr{A}_{0}(p,\mathbf{x})$ =$=\{(g_{1},\cdots,g_{n})\in(\Gamma\mathcal{G})^{n}\mid\forall\,i\in\{1,\cdots,n\},\,p_{[x_{i-1},x_{i}]}\subset g_{1}\cdots g_{i}.\Gamma\},$
$\mathscr{A}_{1}(p,\mathbf{x})$ =$=\{(g_{1},\cdots,g_{n})\in\mathscr{A}_{0}(p,\mathbf{x})\mid\,\|g_{1}\cdots g_{n}\|=\|g_{1}\|+\cdots+\|g_{n}\|\,\}.$
@@
    If there exists $v\in\mathrm{St}_{\Gamma}(v(x_{i-1}))\cap\mathrm{St}_{\Gamma}(v(x_{i+1}))$ with $i\in\{1,\cdots,n-1\}$ such that $\mathrm{supp}(g_{i+1})\subset\mathrm{Lk}_{\Gamma}(v)$, then $\{v(x_{i-1}),v(x_{i+1})\}\subset\mathrm{Lk}_{\Gamma}(v(x_{i}))$ and for any $h\in G_{v}$, we have $(g_{1},\cdots,g_{i}h,g_{i+1},h^{-1}g_{i+2},\cdots,g_{n})\in\mathscr{A}_{0}(p,\mathbf{x})$, where this means $(g_{1},\cdots,g_{n-1}h,g_{n})\in\mathscr{A}_{0}(p,\mathbf{x})$ when $i=n-1$.
@@
    If there exists $v\in\mathrm{St}_{\Gamma}(v(x_{i-1}))\cap\mathrm{St}_{\Gamma}(v(x_{j-1}))$ with $11\leq i<j\leq n+$ 1 such that $\bigcup_{i<k<j}\mathrm{supp}(g_{k})\subset\mathrm{Lk}_{\Gamma}(v)$, then $j\leq i+2$.
@@ 3.4 Classification of geodesic bigons and triangles
When $12\leq i\leq n-$ 1, we have $V(p_{[x_{i-1},x_{i}]})\cap V(q_{[y_{i-1},y_{i}]})=\emptyset$. Since we have $d_{\Gamma^{e}}(y_{i-1},y_{i})\leq d_{\Gamma^{e}}(y_{i-1},x_{i-1})+d_{\Gamma^{e}}(x_{i-1},x_{i})+d_{\Gamma^{e}}(x_{i},y_{i})\leq 10$, the subpath $q_{[y_{i-1},y_{i}]}$ is a unique geodesic in $\Gamma^{e}$ form $y_{i-1}$ to $y_{i}$ by $\mathrm{girth}(\Gamma)>20$. Hence, one of (i) - (iv) below holds, (i) $y_{i-1}=y_{i}$, (ii) $x_{i-1}=x_{i}$, (iii) the sequence $(x_{i-1},x_{i},y_{i-1},y_{i})$ is a geodesic path in $g_{i}.\Gamma$, (iv) the sequence $(y_{i-1},y_{i},x_{i-1},x_{i})$ is a geodesic path in $g_{i}.\Gamma$.
@@
Let $a,b,c\in V(\Gamma^{e})$ be distinct and let $p\in\mathrm{Geo}_{\Gamma^{e}}(a,b)$, $q\in\mathrm{Geo}_{\Gamma^{e}}(b,c)$, and $r\in\mathrm{Geo}_{\Gamma^{e}}(a,c)$ such that the triangle $pqr^{-1}$ is a circuit. Then, there exist a subsequence $(a\!=)\,x_{0},\cdots,x_{n},x^{\prime}_{0},\cdots,x^{\prime}_{m}\,(=\!b)$ of $V(p)$, a subsequence $(b\!=)\,y_{m},\cdots,y_{0},y^{\prime}_{0},\cdots,y^{\prime}_{\ell}\,(=\!c)$ of $V(q)$, a subsequence $(a\!=)\,z_{0},\cdots,z_{n},z^{\prime}_{0},\cdots,z^{\prime}_{\ell}\,(=\!c)$ of $V(r)$, and $f_{1},\cdots,f_{n},g_{1}\cdots,g_{m},h_{1}\cdots,h_{\ell},k\in\Gamma\mathcal{G}$ that satisfy the following 8 conditoins.
@@
When $|\gamma_{[z_{N},z_{N+1}]}|<7$ and $N\geq 1$, we have $d_{\Gamma^{e}}(z_{N},z^{\prime})\leq d_{\Gamma^{e}}(z_{N},z_{N+1})+d_{\Gamma^{e}}(z_{N+1},z^{\prime})\leq 2+7=9$. This implies $\gamma_{[z_{N},b^{\prime}]}\subset\mathcal{N}_{\Gamma^{e}}(b,r+18)$ in the same way as ()($\ast$)( ∗ ). Note $d_{\Gamma^{e}}(z_{N},x_{N})\leq 2$ by $g_{N}\neq g_{N+1}$ and Remark 3.12 (3). Hence, for any $v\in p_{[x_{N},b]}$, we have $d_{\Gamma^{e}}(b,v)\leq d_{\Gamma^{e}}(b,x_{N})\leq d_{\Gamma^{e}}(b,z^{\prime})+d_{\Gamma^{e}}(z^{\prime},z_{N})+d_{\Gamma^{e}}(z_{N},x_{N})\leq r+9+2=r+11$. Thus, (ii) holds by defining $n$ by $n=N$.
When $|\gamma_{[z_{N},z_{N+1}]}|<7$ and $N=0$, we have $d_{\Gamma^{e}}(z_{0},z^{\prime})\leq d_{\Gamma^{e}}(z_{0},z_{1})+d_{\Gamma^{e}}(z_{1},z^{\prime})\leq 9$, hence (i) holds by the same computation as ()($\ast$)( ∗ ).
@@
When $m-1>n-k$, we can also see $w_{m-(n-k)-1}\in\gamma_{[z_{n-1},b^{\prime}]}$. If $w_{m-(n-k)-1}\in\gamma_{[z_{n-1},z_{n}]}$, then we have $|\gamma_{[w_{m-(n-k)-1},z_{n}]}|\leq 2$ by $|\gamma_{[w_{m-(n-k)-1},w_{m-(n-k)-2}]}|\geq 7$ and Remark 3.12 (3). Indeed, $|\gamma_{[w_{m-(n-k)-1},z_{n}]}|\geq 3$ would imply $h_{m-(n-k)-1}=g_{n}(=h_{m-(n-k)})$, which contradicts $h_{m-(n-k)-1}\neq h_{m-(n-k)}$. Hence, we have $d_{\Gamma^{e}}(w_{m-(n-k)-1},b)\leq(r+18)+2$ by $\gamma_{[z_{n},b^{\prime}]}\subset\mathcal{N}_{\Gamma^{e}}(b,r+18)$ in (ii). This and $d_{\Gamma^{e}}(y_{m-(n-k)-1},w_{m-(n-k)-1})\leq 2$ imply $d_{\Gamma^{e}}(y_{m-(n-k)-1},b)\leq d_{\Gamma^{e}}(y_{m-(n-k)-1},w_{m-(n-k)-1})+d_{\Gamma^{e}}(y_{m-(n-k)-1},b)\leq r+22$. Recall $d_{\Gamma^{e}}(b,x_{n})\leq r+18$ in (ii) and $d_{\Gamma^{e}}(a,y_{m})\leq r+11$ in (ii’). Thus, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{m-(n-k)-1})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement.
@@
When $m-1=n-k$, we can also see $w_{0}\in\gamma_{[z_{n-1},b^{\prime}]}$. If $w_{0}\in\gamma_{[z_{n},b^{\prime}]}$, then we have $d_{\Gamma^{e}}(b,y_{0})=d_{\Gamma^{e}}(b,w_{0})\leq r+18$ by (ii), hence the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. If $w_{0}\in\gamma_{[z_{n-1},z_{n}]}$, then we have $z_{n}\in\gamma_{[w_{0},b^{\prime}]}=q_{[w_{0},b^{\prime}]}=q_{[y_{0},b^{\prime}]}$, which implies $|q_{[y_{1},z_{n}]}|\geq|q_{[y_{1},y_{0}]}|\geq 7$ and $p_{[x_{n-1},x_{n}]}\cup q_{[y_{1},z_{n}]}\subset g_{n}.\Gamma$ by $\{y_{1},z_{n}\}\subset h_{1}.\Gamma=g_{n}.\Gamma$ and Corollary 3.30. Hence, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{n})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement.
@@
When $m-1<n-k$, we can also see $w_{0}\in\gamma_{[z_{k+m-2},z_{k+m}]}$ and $w_{1}\in\gamma_{[a,z_{k+m-1}]}$. If $w_{0}\in\gamma_{[z_{k+m-2},z_{k+m-1}]}$, then we have $z_{k+m-1},\cdots,z_{n}\in\gamma_{[w_{0},b^{\prime}]}=q_{[w_{0},b^{\prime}]}=q_{[y_{0},b^{\prime}]}$, which implies $|q_{[y_{1},z_{k+m-1}]}|\geq|q_{[y_{1},y_{0}]}|\geq 7$ and $p_{[x_{k+m-2},x_{k+m-1}]}\cup q_{[y_{1},z_{k+m-1}]}\subset g_{k+m-1}.\Gamma$ by $\{y_{1},z_{k+m-1}\}\subset h_{1}.\Gamma=g_{k+m-1}.\Gamma$ and Corollary 3.30. Hence, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{k+m-1},\cdots,z_{n})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. If $w_{0}\in\gamma_{[z_{k+m-1},z_{k+m}]}$, then by $w_{1}\in\gamma_{[a,z_{k+m-1}]}$, $g_{k+m}\neq g_{k+m-1}=h_{1}$, and Remark 3.12 (3), we have $|\gamma_{[z_{k+m-1},y_{0}]}|=|\gamma_{[z_{k+m-1},w_{0}]}|\leq 2$. We also have $z_{k+m},\cdots,z_{n}\in\gamma_{[w_{0},b^{\prime}]}=q_{[w_{0},b^{\prime}]}=q_{[y_{0},b^{\prime}]}$. Hence, $|q_{[y_{0},z_{k+m}]}|=|\gamma_{[y_{0},z_{k+m}]}|=|\gamma_{[z_{k+m-1},z_{k+m}]}|-|\gamma_{[z_{k+m-1},w_{0}]}|\geq 7-2=5$. Hence, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0},z_{k+m},\cdots,z_{n})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement.
@@
If $h_{i}=g_{1}$, then we have $p_{[w_{i},x_{1}]}\cup q_{[y_{i},y_{i-1}]}\subset h_{i}.\Gamma$ by $\{w_{i},x_{1}\}\subset h_{i}.\Gamma$ and Corollary 3.30. Note that we have $|p_{[w_{i},x_{1}]}|\geq|p_{[x_{0},x_{1}]}|\geq 7$ when $n>1$. In the same way as case (A2), we can show $h_{i-(j-1)}=g_{j}$ for any $1\leq j\leq\min\{i,n\}$. Hence, when either $i>n$ or $[i=n~{}\wedge~{}y_{0}=w_{0}\in\gamma_{[z_{n},b^{\prime}]}]$ holds, the sequences $(w_{m},\cdots,w_{i},x_{1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{i-n})\subset V(q)$, and $(h_{m},\cdots,h_{i-(n-1)})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i\leq n$ and $y_{0}=w_{0}\in\gamma_{[z_{i-1},z_{i}]})$, the sequences $(w_{m},\cdots,w_{i},x_{1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{i},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i+1},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i<n$ and $y_{0}=w_{0}\in\gamma_{[z_{i},z_{i+1}]}$, the sequences $(w_{m},\cdots,w_{i},x_{1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0},z_{i+1},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i+1},\cdots,n)\subset\Gamma\mathcal{G}$ satisfy the statement.
If $h_{i}\neq g_{1}$, then we have $|\gamma_{[z_{0},w_{i-1}]}|\leq 2$ by Remark 3.12 (3). Hence, $|\gamma_{[w_{i},x_{0}]}|=|\gamma_{[w_{i},z_{0}]}|=|\gamma_{[w_{i},w_{i-1}]}|-|\gamma_{[z_{0},w_{i-1}]}|\geq 7-2=5$. Note $p_{[w_{i},x_{0}]}\cup q_{[y_{i},y_{i-1}]}\subset h_{i}.\Gamma$. In the same way as case (A2), when $i>1$, we can show $h_{i-j}=g_{j}$ for any $1\leq j\leq\min\{i-1,n\}$. Hence, when either $i-1>n$ or $[i-1=n~{}\wedge~{}y_{0}=w_{0}\in\gamma_{[z_{n},b^{\prime}]}]$ holds, the sequences $(w_{m},\cdots,w_{i},x_{0},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{i-n-1})\subset V(q)$, and $(h_{m},\cdots,h_{i-n})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i-1\leq n$ and $y_{0}=w_{0}\in\gamma_{[z_{i-2},z_{i-1}]})$, the sequences $(w_{m},\cdots,w_{i},x_{0},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{i-1},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i-1<n$ and $y_{0}=w_{0}\in\gamma_{[z_{i-1},z_{i}]}$, the sequences $(w_{m},\cdots,w_{i},x_{0},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0},z_{i},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. ∎

### v3

@@ 3.2 Connection to the crossing graph and the contact graph
$E(\widehat{\Gamma}^{e})$ $=\{(x,y)\in V(\Gamma^{e})\times V(\Gamma^{e})\mid\text{ $x\neq y$ and $\exists\,g\in\Gamma\mathcal{G},\{x,y\}\subset g.\Gamma$}\}$
@@
Suppose $(g.v,h.w)\in E(\Gamma^{e})$, then $(v,w)\in E(\Gamma)$ and $(g.v,h.w)=(k.v,k.w)$ for some $k\in\Gamma\mathcal{G}$ by Lemma 3.8. This implies $F(g.v)=kJ_{v}$ and $F(h.w)=kJ_{w}$. By taking $a\in G_{v}\setminus\{1\}$ and $b\in G_{w}\setminus\{1\}$, we get the square induced by the vertices $\{k,ka,kb,kab\}$ in $X$. Hence, $F(g.v)$ and $F(h.w)$ are adjacent in $\Delta X$. Conversely, suppose that $gJ_{v}$ and $hJ_{w}$ are adjacent in $\Delta X$. By $N^{\prime}(gJ_{v})\cap N^{\prime}(hJ_{w})\neq\emptyset$ and Proposition 2.21 (3), there exist $s,t\in\Gamma\mathcal{G}$ with $\mathrm{supp}(s)\subset\mathrm{St}_{\Gamma}(v)$ and $\mathrm{supp}(t)\subset\mathrm{St}_{\Gamma}(w)$ such that $gs=ht$. This implies $g.v=gs.v$ and $h.w=g(g^{-1}h).w=g(st^{-1}).w=gs.w$. Also, $(v,w)\in E(\Gamma)$ by Proposition 2.21 (4). Hence, $(g.v,h.w)=gs.(v,w)\in E(\Gamma^{e})$. Thus, $F$ is an isomorphism from $\Gamma^{e}$ to $\Delta X$.
Suppose $(g.v,h.w)\in E(\widehat{\Gamma}^{e})$, then $(g.v,h.w)=(k.v,k.w)$ for some $k\in\Gamma\mathcal{G}$. This implies $k\in N^{\prime}(gJ_{v})\cap N^{\prime}(hJ_{w})$. Hence, $gJ_{v}$ and $hJ_{w}$ are adjacent in $\mathcal{C}X$. The converse direction can be shown in the same way as the argument for $\Gamma^{e}$. Thus, $F\colon\widehat{\Gamma}^{e}\to\mathcal{C}X$ is isomorphic. ∎
@@ 3.3 Admissible paths
$\mathscr{A}_{0}(p,\mathbf{x})$ $=\{(g_{1},\cdots,g_{n})\in(\Gamma\mathcal{G})^{n}\mid\forall\,i\in\{1,\cdots,n\},\,p_{[x_{i-1},x_{i}]}\subset g_{1}\cdots g_{i}.\Gamma\},$
$\mathscr{A}_{1}(p,\mathbf{x})$ $=\{(g_{1},\cdots,g_{n})\in\mathscr{A}_{0}(p,\mathbf{x})\mid\,\|g_{1}\cdots g_{n}\|=\|g_{1}\|+\cdots+\|g_{n}\|\,\}.$
@@
    If there exists $v\in\mathrm{St}_{\Gamma}(v(x_{i-1}))\cap\mathrm{St}_{\Gamma}(v(x_{i+1}))$ with $i\in\{1,\cdots,n-1\}$ such that $\mathrm{supp}(g_{i+1})\subset\mathrm{Lk}_{\Gamma}(v)$, then $\{v(x_{i-1}),v(x_{i+1})\}\subset\mathrm{Lk}_{\Gamma}(v(x_{i}))$ and for any $h\in G_{v}$, we have $(g_{1},\cdots,g_{i}h,g_{i+1},h^{-1}g_{i+2},\cdots,g_{n})\in\mathscr{A}_{0}(p,\mathbf{x})$, where this means $(g_{1},\cdots,g_{n-1}h,g_{n})\in\mathscr{A}_{0}(p,\mathbf{x})$ when i=n1i=n-1.
@@
    If there exists $v\in\mathrm{St}_{\Gamma}(v(x_{i-1}))\cap\mathrm{St}_{\Gamma}(v(x_{j-1}))$ with $1\leq i<j\leq n+1$ such that $\bigcup_{i<k<j}\mathrm{supp}(g_{k})\subset\mathrm{Lk}_{\Gamma}(v)$, then $j\leq i+2$.
@@ 3.4 Classification of geodesic bigons and triangles
When $2\leq i\leq n-1$, we have $V(p_{[x_{i-1},x_{i}]})\cap V(q_{[y_{i-1},y_{i}]})=\emptyset$. Since we have $d_{\Gamma^{e}}(y_{i-1},y_{i})\leq d_{\Gamma^{e}}(y_{i-1},x_{i-1})+d_{\Gamma^{e}}(x_{i-1},x_{i})+d_{\Gamma^{e}}(x_{i},y_{i})\leq 10$, the subpath $q_{[y_{i-1},y_{i}]}$ is a unique geodesic in $\Gamma^{e}$ form $y_{i-1}$ to $y_{i}$ by $\mathrm{girth}(\Gamma)>20$. Hence, one of (i) - (iv) below holds, (i) $y_{i-1}=y_{i}$, (ii) $x_{i-1}=x_{i}$, (iii) the sequence $(x_{i-1},x_{i},y_{i-1},y_{i})$ is a geodesic path in $g_{i}.\Gamma$, (iv) the sequence $(y_{i-1},y_{i},x_{i-1},x_{i})$ is a geodesic path in $g_{i}.\Gamma$.
@@
Let $a,b,c\in V(\Gamma^{e})$ be distinct and let $p\in\mathrm{Geo}_{\Gamma^{e}}(a,b)$, $q\in\mathrm{Geo}_{\Gamma^{e}}(b,c)$, and $r\in\mathrm{Geo}_{\Gamma^{e}}(a,c)$ such that the triangle $pqr^{-1}$ is a circuit. Then, there exist a subsequence $(a\!=)\,x_{0},\cdots,x_{n},x^{\prime}_{0},\cdots,x^{\prime}_{m}\,(=\!b)$ of $V(p)$, a subsequence $(b\!=)\,y_{m},\cdots,y_{0},y^{\prime}_{0},\cdots,y^{\prime}_{\ell}\,(=\!c)$ of $V(q)$, a subsequence $(a\!=)\,z_{0},\cdots,z_{n},z^{\prime}_{0},\cdots,z^{\prime}_{\ell}\,(=\!c)$ of $V(r)$, and $f_{1},\cdots,f_{n},g_{1}\cdots,g_{m},h_{1}\cdots,h_{\ell},k\in\Gamma\mathcal{G}$ that satisfy the following 8 conditions.
@@
When $|\gamma_{[z_{N},z_{N+1}]}|<7$ and $N\geq 1$, we have $d_{\Gamma^{e}}(z_{N},z^{\prime})\leq d_{\Gamma^{e}}(z_{N},z_{N+1})+d_{\Gamma^{e}}(z_{N+1},z^{\prime})\leq 2+7=9$. This implies $\gamma_{[z_{N},b^{\prime}]}\subset\mathcal{N}_{\Gamma^{e}}(b,r+18)$ in the same way as $(\ast)$. Note $d_{\Gamma^{e}}(z_{N},x_{N})\leq 2$ by $g_{N}\neq g_{N+1}$ and Remark 3.12 (3). Hence, for any $v\in p_{[x_{N},b]}$, we have $d_{\Gamma^{e}}(b,v)\leq d_{\Gamma^{e}}(b,x_{N})\leq d_{\Gamma^{e}}(b,z^{\prime})+d_{\Gamma^{e}}(z^{\prime},z_{N})+d_{\Gamma^{e}}(z_{N},x_{N})\leq r+9+2=r+11$. Thus, (ii) holds by defining $n$ by $n=N$.
When $|\gamma_{[z_{N},z_{N+1}]}|<7$ and $N=0$, we have $d_{\Gamma^{e}}(z_{0},z^{\prime})\leq d_{\Gamma^{e}}(z_{0},z_{1})+d_{\Gamma^{e}}(z_{1},z^{\prime})\leq 9$, hence (i) holds by the same computation as $(\ast)$.
@@
When m1>nkm-1>n-k, we can also see $w_{m-(n-k)-1}\in\gamma_{[z_{n-1},b^{\prime}]}$. If $w_{m-(n-k)-1}\in\gamma_{[z_{n-1},z_{n}]}$, then we have $|\gamma_{[w_{m-(n-k)-1},z_{n}]}|\leq 2$ by $|\gamma_{[w_{m-(n-k)-1},w_{m-(n-k)-2}]}|\geq 7$ and Remark 3.12 (3). Indeed, $|\gamma_{[w_{m-(n-k)-1},z_{n}]}|\geq 3$ would imply $h_{m-(n-k)-1}=g_{n}(=h_{m-(n-k)})$, which contradicts $h_{m-(n-k)-1}\neq h_{m-(n-k)}$. Hence, we have $d_{\Gamma^{e}}(w_{m-(n-k)-1},b)\leq(r+18)+2$ by $\gamma_{[z_{n},b^{\prime}]}\subset\mathcal{N}_{\Gamma^{e}}(b,r+18)$ in (ii). This and $d_{\Gamma^{e}}(y_{m-(n-k)-1},w_{m-(n-k)-1})\leq 2$ imply $d_{\Gamma^{e}}(y_{m-(n-k)-1},b)\leq d_{\Gamma^{e}}(y_{m-(n-k)-1},w_{m-(n-k)-1})+d_{\Gamma^{e}}(y_{m-(n-k)-1},b)\leq r+22$. Recall $d_{\Gamma^{e}}(b,x_{n})\leq r+18$ in (ii) and $d_{\Gamma^{e}}(a,y_{m})\leq r+11$ in (ii’). Thus, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{m-(n-k)-1})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement.
@@
When m1=nkm-1=n-k, we can also see $w_{0}\in\gamma_{[z_{n-1},b^{\prime}]}$. If $w_{0}\in\gamma_{[z_{n},b^{\prime}]}$, then we have $d_{\Gamma^{e}}(b,y_{0})=d_{\Gamma^{e}}(b,w_{0})\leq r+18$ by (ii), hence the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. If $w_{0}\in\gamma_{[z_{n-1},z_{n}]}$, then we have $z_{n}\in\gamma_{[w_{0},b^{\prime}]}=q_{[w_{0},b^{\prime}]}=q_{[y_{0},b^{\prime}]}$, which implies $|q_{[y_{1},z_{n}]}|\geq|q_{[y_{1},y_{0}]}|\geq 7$ and $p_{[x_{n-1},x_{n}]}\cup q_{[y_{1},z_{n}]}\subset g_{n}.\Gamma$ by $\{y_{1},z_{n}\}\subset h_{1}.\Gamma=g_{n}.\Gamma$ and Corollary 3.29. Hence, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{n})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement.
@@
When m1<nkm-1<n-k, we can also see $w_{0}\in\gamma_{[z_{k+m-2},z_{k+m}]}$ and $w_{1}\in\gamma_{[a,z_{k+m-1}]}$. If $w_{0}\in\gamma_{[z_{k+m-2},z_{k+m-1}]}$, then we have $z_{k+m-1},\cdots,z_{n}\in\gamma_{[w_{0},b^{\prime}]}=q_{[w_{0},b^{\prime}]}=q_{[y_{0},b^{\prime}]}$, which implies $|q_{[y_{1},z_{k+m-1}]}|\geq|q_{[y_{1},y_{0}]}|\geq 7$ and $p_{[x_{k+m-2},x_{k+m-1}]}\cup q_{[y_{1},z_{k+m-1}]}\subset g_{k+m-1}.\Gamma$ by $\{y_{1},z_{k+m-1}\}\subset h_{1}.\Gamma=g_{k+m-1}.\Gamma$ and Corollary 3.29. Hence, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{k+m-1},\cdots,z_{n})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. If $w_{0}\in\gamma_{[z_{k+m-1},z_{k+m}]}$, then by $w_{1}\in\gamma_{[a,z_{k+m-1}]}$, $g_{k+m}\neq g_{k+m-1}=h_{1}$, and Remark 3.12 (3), we have $|\gamma_{[z_{k+m-1},y_{0}]}|=|\gamma_{[z_{k+m-1},w_{0}]}|\leq 2$. We also have $z_{k+m},\cdots,z_{n}\in\gamma_{[w_{0},b^{\prime}]}=q_{[w_{0},b^{\prime}]}=q_{[y_{0},b^{\prime}]}$. Hence, $|q_{[y_{0},z_{k+m}]}|=|\gamma_{[y_{0},z_{k+m}]}|=|\gamma_{[z_{k+m-1},z_{k+m}]}|-|\gamma_{[z_{k+m-1},w_{0}]}|\geq 7-2=5$. Hence, the sequences $(x_{k-1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0},z_{k+m},\cdots,z_{n})\subset V(q)$, and $(g_{k},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement.
@@
If $h_{i}=g_{1}$, then we have $p_{[w_{i},x_{1}]}\cup q_{[y_{i},y_{i-1}]}\subset h_{i}.\Gamma$ by $\{w_{i},x_{1}\}\subset h_{i}.\Gamma$ and Corollary 3.29. Note that we have $|p_{[w_{i},x_{1}]}|\geq|p_{[x_{0},x_{1}]}|\geq 7$ when $n>1$. In the same way as case (A2), we can show $h_{i-(j-1)}=g_{j}$ for any $1\leq j\leq\min\{i,n\}$. Hence, when either $i>n$ or $i=n~\wedge~y_{0}=w_{0}\in\gamma_{[z_{n},b^{\prime}]}]$ holds, the sequences $(w_{m},\cdots,w_{i},x_{1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{i-n})\subset V(q)$, and $(h_{m},\cdots,h_{i-(n-1)})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i\leq n$ and $y_{0}=w_{0}\in\gamma_{[z_{i-1},z_{i}]})$, the sequences $(w_{m},\cdots,w_{i},x_{1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{i},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i+1},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i<n$ and $y_{0}=w_{0}\in\gamma_{[z_{i},z_{i+1}]}$, the sequences $(w_{m},\cdots,w_{i},x_{1},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0},z_{i+1},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i+1},\cdots,n)\subset\Gamma\mathcal{G}$ satisfy the statement.
If $h_{i}\neq g_{1}$, then we have $|\gamma_{[z_{0},w_{i-1}]}|\leq 2$ by Remark 3.12 (3). Hence, $|\gamma_{[w_{i},x_{0}]}|=|\gamma_{[w_{i},z_{0}]}|=|\gamma_{[w_{i},w_{i-1}]}|-|\gamma_{[z_{0},w_{i-1}]}|\geq 7-2=5$. Note $p_{[w_{i},x_{0}]}\cup q_{[y_{i},y_{i-1}]}\subset h_{i}.\Gamma$. In the same way as case (A2), when $i>1$, we can show $h_{i-j}=g_{j}$ for any $1\leq j\leq\min\{i-1,n\}$. Hence, when either i1>ni-1>n or $i-1=n~\wedge~y_{0}=w_{0}\in\gamma_{[z_{n},b^{\prime}]}]$ holds, the sequences $(w_{m},\cdots,w_{i},x_{0},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{i-n-1})\subset V(q)$, and $(h_{m},\cdots,h_{i-n})\subset\Gamma\mathcal{G}$ satisfy the statement. When $i-1\leq n$ and $y_{0}=w_{0}\in\gamma_{[z_{i-2},z_{i-1}]})$, the sequences $(w_{m},\cdots,w_{i},x_{0},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{1},z_{i-1},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. When i1<ni-1<n and $y_{0}=w_{0}\in\gamma_{[z_{i-1},z_{i}]}$, the sequences $(w_{m},\cdots,w_{i},x_{0},\cdots,x_{n})\subset V(p)$, $(y_{m},\cdots,y_{0},z_{i},\cdots,z_{n})\subset V(q)$, and $(h_{m},\cdots,h_{1},g_{i},\cdots,g_{n})\subset\Gamma\mathcal{G}$ satisfy the statement. ∎